# Remarks on Dirac–Bergmann algorithm, Dirac’s conjecture and the extended Hamiltonian

### v2

@@ 3.2 On the extended Hamiltonian
Why constrained modes matter. Interaction with constrained modes. One might object to our discussion by noting that constrained modes are often treated as redundant, unphysical variables, and it is commonly stated that only dynamical modes carry physical content. If we solve the constraint equation in (7) for $y$ in terms of $\dot{x}$ and substitute back into the Lagrangian, we obtain $L=0$, i.e. an empty system. If both Hamiltonian descriptions have no dynamical modes, then why should the additional gauge freedom of the extended Hamiltonian be a problem? The answer is that, in such situations, the notion that constrained modes are unphysical is not a useful one. Expressing constrained variables through the other ones does not remove the associated physical information. It is simply transferred into the remaining equations of motion and becomes manifest in the presence of couplings. In gauge theories this information encodes the interaction of static configurations of gauge fields with matter fields.
@@
The first two equations imply $\frac{d}{dt}(q^{\ast}q)=0$, i.e. conservation of the charge $q^{\ast}q=C$. If we again solve the second equation for $y$ and substitute it in the Lagrangian, we obtain $L=q^{\ast}(i\>\partial_{t}+\dot{x})q+\frac{1}{2}(q^{\ast}q)^{2}$. Varying the action then yields the same charge conservation and the equation $(i\>\partial_{t}+\dot{x}+q^{\ast}q)q=0$. The residual gauge symmetry allows us to set $x=0$. Using charge conservation, the equation reduces to $(i\>\partial_{t}+C)q=0$. Thus the presence of interaction changes the equations for $q$, even though the $(x,y)$-sector still has no dynamical modes.
When one passes to the extended Hamiltonian, the equation is changed to $(i\>\partial_{t}+C+\lambda_{x})q=0$. The charge remains conserved, but the arbitrary Lagrange multiplier $\lambda_{x}$ allows one to absorb its contribution. Now the system can be interpreted as a $L=q^{\ast}(i\>\partial_{t})q$, i.e. the $q$-sector without interaction.
@@
The structure of the new system, however, suggests a natural way to resolve this mismatch. One must reinterpret the variables. For our toy model, the guiding principle is straightforward: the physically meaningful gauge invariant combination in the Lagrangian picture, $\dot{x}+y=0$, should be identified with the momentum $p_{x}=0$, as it is gauge invariant under transformations generated by the Dirac generator $G_{D}=(-\dot{\varepsilon}+\xi)\>p_{y}+\varepsilon\>p_{y}$. The price is that the set of solutions for the variables $x(t)$ and $y(t)$ in the extended Hamiltonian framework is larger than in the original formulation. This, however, is not problematic as gauge-noninvariant quantities are not observable in any case.
The interpretation of this type was adopted in [24], where Dirac’s conjecture was claimed to be proven for an arbitrary canonical gauge theory. The key step in their proof was, in fact, to redefine the system so that the conjecture holds: one demands that all first-class constraints generate gauge transformations and, correspondingly, declares as gauge invariant those phase-space functions that commute with all first-class constraints on the constraint surface. Such quantities are usually called canonical gauge invariants [^13] Canonical gauge invariant quantities are not, in general, the same as gauge invariants in the Lagrangian formalism. For example, the left-hand side of Gauß’s law in Yang–Mills theory $(D_{i}F^{i0})_{a}=0$ is not a scalar with respect to the internal symmetry group, but it is a canonical gauge invariant quantity if expressed through canonical variables..
With this definition, the set of canonical gauge invariants is the same in both the total Hamiltonian and the extended Hamiltonian approaches. Indeed, a phase-space function can be invariant under the action of (15) only if it commutes with all first-class constraints, since $\varepsilon$ is arbitrary. One can further easily show that the dynamics of these invariants agrees with that obtained in the total Hamiltonian approach [24, 18]. In electrodynamics, this redefinition means identifying the electric field strength with the spatial momenta and abandoning its standard definition in the longitudinal sector [17, 22].
To see why such a redefinition works, it is useful to view it from a different angle. Let us return to the system with the total Hamiltonian (9). In this formulation, the gauge symmetry is generated by (17).
As we have mentioned, the equations (10) can be obtained from the first-order Lagrangian
@@
This substitution “splits” the original variable $y$ into two components and thereby effectively introduces an additional gauge redundancy. The variables $y^{\scriptscriptstyle E}$ and $\lambda_{x}$ can now be transformed independently. After the substitution, we arrive at
@@
where $\lambda^{\scriptscriptstyle E}_{y}=\lambda_{y}+\dot{\lambda}^{\scriptscriptstyle E}_{x}$. The first-order Lagrangian (24) reproduces precisely the equations of the extended Hamiltonian system (19). Equivalently, one may carry out the substitution directly in the equations (10), which leads to the very same result. The redefined system is equivalent to $L=\frac{1}{2}(\dot{x}+y^{\scriptscriptstyle E}-\lambda_{x})^{2}$. In fact, the procedure we have applied is nothing but the Stückelberg trick. The basic idea is to introduce new variables by reparameterizing the original ones, thereby adding an extra gauge redundancy. A familiar example is an artificial introduction of $U(1)$ gauge invariance for a massive vector field, achieved in essentially the same manner [17].
The constrained mode is saved – it is now encoded in the momentum $p_{x}$. This can be formalized as follows. Using our previously introduced prescription for the number of modes, we see that the physical-sector result coincides with that obtained in the total Hamiltonian approach, namely $N_{dyn}=0$, $N_{con}=1$, $N_{gauge}=2$, provided that the new Lagrange multiplier is included in the total number of variables, so that $N_{tot}$ increases to $3$. This is not surprising: we have deliberately enlarged the space of variables by adding an additional pure-gauge degree of freedom $\lambda_{x}$, which, as shown above, can be introduced already at the Lagrangian level without changing the physical content.
@@ 4 Hamiltonian formalism of common gauge field theories
From now on, we will always implicitly assume that modes are counted in the local sense, i.e. in the presence of appropriate fall-off boundary conditions for the fields.
4.1 On canonical formulation of Yang–Mills theory
Let us begin by considering Yang–Mills theory with an arbitrary $N$-dimensional simple gauge group without sources. The Lagrangian density of Yang–Mills theory is given by
@@ 4.1 On canonical formulation of Yang–Mills theory
where $F^{a}_{\mu\nu}=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}+g\>C^{abc}A^{b}_{\mu}A^{c}_{\nu}$ is the field strength tensor. $C^{abc}$ are the structure constants. The equations of motion of Yang–Mills theory have the form
@@
As expected, the equation for the time derivative of $A^{a}_{i}$ reproduces the definition of the momenta (29). Substituting this relation into the remaining equations for the momenta yields the Yang–Mills equations of motion (27).
@@ 4.1.1 Generators of gauge transformations
Again, the Dirac conjecture does not hold for the Hamiltonian formulation of Yang–Mills theory in the standard sense of gauge transformations. To see this, it is sufficient to demonstrate that the Castellani generator produces the gauge group transformations (28), whereas the Dirac generator does not.
The key difference from the toy model’s generator (17) is that the part of the generator corresponding to a zero-order term is not precisely the secondary constraint. It differs from it by a combination of constraints which one has to add in order to compensate the contribution $\{(D_{i}\pi^{i})^{a},H_{T}\}=-g\>C^{abc}A^{b}_{0}(D_{j}\pi^{j})^{c}$ in the chain:
@@
The temporal and spatial components of the Yang–Mills field can now be transformed independently of one another, which is a broader symmetry than that of the original system. This, however, is correct only when the physical state is considered at a fixed instant of time in the total Hamiltonian approach.
@@
The logic is the same as before. If we now include the $N$ new variables $\lambda^{a}_{1}$ in the total number of modes $N_{tot}$, we obtain $N_{dyn}=2N,\ N_{gauge}=2N,\ N_{con}=N.$ Thus the number of physical degrees of freedom remains unchanged, while the number of gauge modes increases by a factor of two. In this sense Dirac’s conjecture becomes valid in both interpretations of symmetry transformation. The gauge transformations now take the form (35).
@@
An analogous situation occurs in electrodynamics [22].
One may object that the chromoelectric field is not, in any case, an observable, since it is not gauge invariant, and therefore changing its definition should not matter. However, one can construct gauge invariant quantities from it, such as $\text{Tr}\>F^{a}_{0i}{}^{2}$, and their relation to the potentials is modified accordingly. More generally, as already noted, what changes is the correspondence between canonical gauge invariant quantities and the variables of the Lagrangian description.
@@
$(D_{i}F_{0i})^{a}=-(D_{i}D_{i}\lambda_{1})^{a},\qquad(D_{0}F_{0i})^{a}-(D_{k}F_{ki})^{a}=-D_{0}(D_{i}\lambda_{1})^{a}$ (39)
which implies that a gauge transformation can generate matter fields out of the vacuum. Exactly the same situation arises in electrodynamics [16]. Thus, we arrive at a modified Yang–Mills theory rather than an equivalent reformulation. In this modified theory, the additional gauge freedom beyond that of the original system is tied to the freedom to choose an arbitrary configuration of effective charges. The constrained modes become spurious, while the number of dynamical modes remains unchanged – we lose the covariantly longitudinal component of the field strength.
@@ 4.2 On the ADM formalism of General Relativity
A crucial difference between general relativity and gauge theories such as Yang–Mills theory is the following: diffeomorphisms, unlike internal symmetry transformations, change not only the functional form of the fields but also the arguments of the fields. Nevertheless, diffeomorphism invariance can be treated formally as an internal symmetry if one considers only the change in the functional dependence of fields. We are therefore interested in the variation of the form $\overline{\delta}g_{\mu\nu}=g^{\prime}_{\mu\nu}(x^{\prime})-g_{\mu\nu}(x)$, which captures how the metric tensor changes as a function of the local coordinates, rather than how its value changes at a fixed point of the manifold. This is nothing but the Lie derivative of the tensor field, actually $\overline{\delta}g_{\mu\nu}=-\mathcal{L}_{\xi}g_{\mu\nu}.$
@@
From the standard transformation law for the metric, we obtain the corresponding expression for the variation of the metric:
@@
Let us consider the gravitational field in the absence of matter. As the action of GR, we take the Einstein–Hilbert action [^17] We are using the signature (,+,+,+)(-,+,+,+).
@@
where $\gamma^{ij}$ is the inverse of the spatial metric $\gamma_{ij}$. The functions $N\equiv N_{0}$ and $N_{i}$ are called the lapse and shift functions, respectively. Equivalently, (44) can be expressed through the metric components as $\gamma_{ij}=g_{ij}$, $g_{0i}=N_{i}$, and $g_{00}=-N^{2}+N_{i}N^{i}$. One can then verify that the inverse metric is given by $g^{00}=-\frac{1}{N^{2}}$, $g^{0i}=\frac{N^{i}}{N^{2}}$, and $g^{ij}=\gamma^{ij}+\frac{g^{0i}g^{0j}}{N^{2}}$ (we assume $g^{00}\neq 0$). Finally, from the expansion of the determinant by algebraic complements it follows that $\sqrt{-g}=N\>\sqrt{\gamma}$.
@@
From the spatial metric $\gamma_{ij}$, one can construct purely three-dimensional geometric objects. We will denote such three-dimensional quantities by placing a superscript $3$ above the symbol, in order to distinguish them from their four-dimensional counterparts; for instance, $\overset{3}{R}_{ijkl}$, $\overset{3}{R}_{ij}$, and $\overset{3}{R}$.
@@
$K_{ij}=\frac{1}{2N}\biggl(\overset{3}{\nabla}_{i}N_{j}+\overset{3}{\nabla}_{j}N_{i}-\dot{\gamma}_{ij}\biggr),\qquad R_{iklm}=\overset{3}{R}_{iklm}-K_{im}K_{kl}+K_{il}K_{km}.$ (45)
@@ 4.2.1 The total Hamiltonian approach
$\mathcal{J}_{ij,kl}$ is the Wheeler–DeWitt metric. It is given by $\mathcal{J}_{ij,kl}=\frac{2\varkappa}{\sqrt{\gamma}}(\gamma_{ik}\gamma_{jl}+\gamma_{il}\gamma_{jk}-\gamma_{ij}\gamma_{kl})$. Its inverse has the form $\mathcal{\hat{J}}^{ij,kl}=\frac{1}{4}\frac{\sqrt{\gamma}}{2\varkappa}(\gamma^{ik}\gamma^{jl}+\gamma^{il}\gamma^{jk}-\gamma^{ij}\gamma^{kl})$.
@@
One can check that the secondary constraints, when expressed through the metric components, are equivalent to the constraints in the Lagrangian formalism $G^{0\mu}=0$, since

### v3

@@ 3.2 On the extended Hamiltonian
Why constrained modes matter. Interaction with constrained modes. One might object to our discussion by noting that constrained modes are often treated as redundant, unphysical variables, and it is commonly stated that only dynamical modes carry physical content. If we solve the constraint equation in (7) for $y$ in terms of $\dot{x}$ and substitute it back into the Lagrangian, we obtain $L=0$, i.e., an empty system. If both Hamiltonian descriptions have no dynamical modes, then why should the additional gauge freedom of the extended Hamiltonian be a problem? The answer is that, in such situations, the notion that constrained modes are unphysical is not a useful one. Once again, expressing constrained variables through the other ones does not remove the associated physical information. It is simply transferred into the remaining equations of motion and becomes manifest in the presence of couplings. In gauge theories this information encodes the interaction of static configurations of gauge fields with matter fields.
@@
The first two equations imply $\frac{d}{dt}(q^{\ast}q)=0$, i.e., conservation of the charge $q^{\ast}q=C$. If we again solve the second equation for $y$ and substitute it in the Lagrangian, we obtain $L=q^{\ast}(i\>\partial_{t}+\dot{x})q+\frac{1}{2}(q^{\ast}q)^{2}$. Varying the action then yields the same charge conservation and the equation $(i\>\partial_{t}+\dot{x}+q^{\ast}q)q=0$. The residual gauge symmetry allows us to set $x=0$. Using charge conservation, the equation reduces to $(i\>\partial_{t}+C)q=0$. Thus the presence of interaction changes the equations for $q$, even though the $(x,y)$-sector still has no dynamical modes.
When one passes to the extended Hamiltonian, the equation is changed to $(i\>\partial_{t}+C+\lambda_{x})q=0$. The charge remains conserved, but the arbitrary Lagrange multiplier $\lambda_{x}$ allows one to absorb its contribution. Now the system can be interpreted as a $L=q^{\ast}(i\>\partial_{t})q$, i.e., the $q$-sector without interaction.
@@
The structure of the new system, however, suggests a natural way to resolve this mismatch. One must reinterpret the variables. For our toy model, the guiding principle is straightforward: the physically meaningful gauge invariant combination in the Lagrangian picture, $\dot{x}+y=0$, should be identified with the momentum $p_{x}=0$, as it is gauge invariant under transformations generated by the Dirac generator $G_{D}=(-\dot{\varepsilon}+\xi)\>p_{y}+\varepsilon\>p_{y}$. The price is that the set of solutions for the variables $x(t)$ and $y(t)$ in the extended Hamiltonian framework is larger than in the original formulation. This, however, is not problematic, as gauge noninvariant quantities are not observable in any case.
The interpretation of this type was adopted in [26], where Dirac’s conjecture was claimed to be proven for an arbitrary canonical gauge theory. The key step in their proof was, in fact, to redefine the system so that the conjecture holds: one demands that all first-class constraints generate gauge transformations and, correspondingly, declares as gauge invariant those phase-space functions that commute with all first-class constraints on the constraint surface. Such quantities are usually called canonical gauge invariants [^13] Canonical gauge invariant quantities are not, in general, the same as gauge invariants in the Lagrangian formalism. For example, the left-hand side of Gauß’s law in Yang-Mills theory $(D_{i}F^{i0})_{a}=0$ is not a scalar with respect to the internal symmetry group, but it is a canonical gauge invariant quantity if expressed through canonical variables (see Section 4)..
With this definition, the set of canonical gauge invariants is the same in both the total Hamiltonian and the extended Hamiltonian approaches. Indeed, a phase-space function can be invariant under the action of (15) only if it commutes with all first-class constraints, since $\varepsilon$ is arbitrary. One can further show that the dynamics of these invariants agree with that obtained in the total Hamiltonian approach [26, 18]. In electrodynamics, this redefinition means identifying the electric field strength with the spatial momenta and abandoning its standard definition in the longitudinal sector [17, 23].
To see why such a redefinition works, it is useful to view it from a different angle. Let us return to the system with the total Hamiltonian (9). In this formulation, the gauge symmetry is generated by (17). As we have mentioned, the equations (10) can be obtained from the first-order Lagrangian
@@
This substitution “splits” the original variable $y$ into two components and thereby effectively introduces an additional gauge redundancy: the variables $y^{\scriptscriptstyle E}$ and $\lambda_{x}$ can now be transformed independently. After the substitution, we arrive at
@@
where $\lambda^{\scriptscriptstyle E}_{y}=\lambda_{y}+\dot{\lambda}^{\scriptscriptstyle E}_{x}$. The first-order Lagrangian (24) reproduces precisely the equations of the extended Hamiltonian system (19). Equivalently, one may carry out the substitution directly in the equations (10). The redefined system is equivalent to $L=\frac{1}{2}(\dot{x}+y^{\scriptscriptstyle E}-\lambda_{x})^{2}$. In fact, the procedure we have applied is nothing but the Stückelberg trick. The basic idea is to introduce new variables by reparameterizing the original ones, thereby adding an extra gauge redundancy. A familiar example is an artificial introduction of $U(1)$ gauge invariance for a massive vector field, achieved in essentially the same manner [17].
The constrained mode is saved, it is now encoded in the momentum $p_{x}$. This can be formalized as follows. Using our previously introduced prescription for the number of modes, we see that the physical-sector result coincides with that obtained in the total Hamiltonian approach, namely $N_{dyn}=0$, $N_{con}=1$, $N_{gauge}=2$, provided that the new Lagrange multiplier is included in the total number of variables, so that $N_{tot}$ increases to $3$. This is not surprising: we have deliberately enlarged the space of variables by adding an additional pure-gauge degree of freedom $\lambda_{x}$, which, as shown above, can be introduced already at the Lagrangian level without changing the physical content.
@@ 4 Hamiltonian formalism of common gauge field theories
From now on, we will always implicitly assume that modes are counted in the local sense, i.e., in the presence of appropriate fall-off boundary conditions for the fields.
4.1 On canonical formulation of Yang-Mills theory
Let us begin by considering Yang-Mills theory with an arbitrary $N$-dimensional simple gauge group without sources. The Lagrangian density of Yang-Mills theory is given by
@@ 4.1 On canonical formulation of Yang-Mills theory
where $F^{a}_{\mu\nu}=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}+g\>C^{abc}A^{b}_{\mu}A^{c}_{\nu}$ is the field strength tensor. $C^{abc}$ are the structure constants. The equations of motion of Yang-Mills theory have the form
@@
As expected, the equation for the time derivative of $A^{a}_{i}$ reproduces the definition of the momenta (29). Substituting this relation into the remaining equations for the momenta yields the Yang-Mills equations of motion (27).
@@ 4.1.1 Generators of gauge transformations
Again, the Dirac conjecture does not hold for the Hamiltonian formulation of Yang-Mills theory in the standard sense of gauge transformations. To see this, it is sufficient to demonstrate that the Castellani generator produces the gauge group transformations (28), whereas the Dirac generator does not.
The key difference from the toy model’s generator (17) is that the part of the generator corresponding to a zero-order term is not precisely the secondary constraint. It differs from it by a combination of constraints which one has to add in order to compensate for the contribution $\{(D_{i}\pi^{i})^{a},H_{T}\}=-g\>C^{abc}A^{b}_{0}(D_{j}\pi^{j})^{c}$ in the chain (16):
@@
The temporal and spatial components of the Yang-Mills field can now be transformed independently of one another, which is a broader symmetry than that of the original system. This, however, is correct in the total Hamiltonian approach only when the physical state is considered at a fixed instant in time.
@@
The logic is the same as before. If we now include the $N$ new variables $\lambda^{a}_{1}$ in the total number of modes $N_{tot}$, we obtain $N_{dyn}=2N,\ N_{gauge}=2N,\ N_{con}=N.$ Thus, the number of physical degrees of freedom remains unchanged, while the number of gauge modes increases by a factor of two. In this sense, Dirac’s conjecture becomes valid in both interpretations of gauge symmetry transformation. The gauge transformations are now given by (35).
@@
One may object that the chromoelectric field is not, in any case, an observable since it is not gauge invariant, and therefore, changing its definition should not matter. However, one can construct gauge invariant quantities from it, such as $\text{Tr}\>F^{a}_{0i}{}^{2}$, and their relation to the potentials is modified accordingly. More generally, as already noted, what changes is the correspondence between canonical gauge invariant quantities and the variables of the Lagrangian description.
@@
$(D_{i}F_{0i})^{a}=-(D_{i}D_{i}\lambda_{1})^{a},\qquad(D_{0}F_{0i})^{a}-(D_{k}F_{ki})^{a}=-D_{0}(D_{i}\lambda_{1})^{a},$ (39)
which implies that a gauge transformation can generate matter fields out of the vacuum. Thus, we arrive at a modified Yang-Mills theory rather than an equivalent reformulation. In this modified theory, the additional gauge freedom beyond that of the original system is tied to the freedom to choose an arbitrary configuration of effective charges. The constrained modes become spurious, while the number of dynamical modes remains unchanged – we lose the covariantly longitudinal component of the chromoelectric field strength.
@@ 4.2 On the ADM formalism of general relativity
A crucial difference between general relativity and gauge theories such as Yang-Mills theory, is the following: diffeomorphisms, unlike internal symmetry transformations, change not only the functional form of the fields but also the arguments of the fields. Nevertheless, diffeomorphism invariance can be treated formally as an internal symmetry if one considers only the change in the functional dependence of fields. We are therefore interested in the variation of the form $\overline{\delta}g_{\mu\nu}=g^{\prime}_{\mu\nu}(x^{\prime})-g_{\mu\nu}(x)$, which captures how the metric tensor changes as a function of the local coordinates, rather than how its value changes at a fixed point of the spacetime manifold. This is nothing but the Lie derivative of the tensor field, actually $\overline{\delta}g_{\mu\nu}=-\mathcal{L}_{\xi}g_{\mu\nu}.$
@@
From the standard transformation law of the metric, we obtain the corresponding expression for its variation:
@@
Let us consider the gravitational field in the absence of matter. The action of GR is the Einstein–Hilbert action [^17] We are using the signature (,+,+,+)(-,+,+,+).
@@
where $\gamma^{ij}$ is the inverse of the spatial metric $\gamma_{ij}$. The functions $N\equiv N_{0}$ and $N_{i}$ are called the lapse and shift functions, respectively. Equivalently (44) can be expressed through the metric components as $\gamma_{ij}=g_{ij}$, $g_{0i}=N_{i}$, and $g_{00}=-N^{2}+N_{i}N^{i}$. One can then verify that the inverse metric is given by $g^{00}=-\frac{1}{N^{2}}$, $g^{0i}=\frac{N^{i}}{N^{2}}$, and $g^{ij}=\gamma^{ij}+\frac{g^{0i}g^{0j}}{N^{2}}$ (we assume $g^{00}\neq 0$). Finally, from the expansion of the determinant by algebraic complements it follows that $\sqrt{-g}=N\>\sqrt{\gamma}$.
@@
From the spatial metric $\gamma_{ij}$, one can construct purely three-dimensional geometric objects. We will denote such three-dimensional quantities by placing a superscript $3$ above the symbol to distinguish them from their four-dimensional counterparts; for instance, $\overset{3}{R}_{ijkl}$, $\overset{3}{R}_{ij}$, and $\overset{3}{R}$.
@@
$K_{ij}=\frac{1}{2N}\biggl(\overset{3}{\nabla}_{i}N_{j}+\overset{3}{\nabla}_{j}N_{i}-\dot{\gamma}_{ij}\biggr)\quad\text{and}\quad R_{iklm}=\overset{3}{R}_{iklm}-K_{im}K_{kl}+K_{il}K_{km}.$ (45)
@@ 4.2.1 The total Hamiltonian approach
$\mathcal{J}_{ij,kl}$ is the Wheeler-DeWitt metric. It is given by $\mathcal{J}_{ij,kl}=\frac{2\varkappa}{\sqrt{\gamma}}(\gamma_{ik}\gamma_{jl}+\gamma_{il}\gamma_{jk}-\gamma_{ij}\gamma_{kl})$. Its inverse has the form $\mathcal{\hat{J}}^{ij,kl}=\frac{1}{4}\frac{\sqrt{\gamma}}{2\varkappa}(\gamma^{ik}\gamma^{jl}+\gamma^{il}\gamma^{jk}-\gamma^{ij}\gamma^{kl})$.
@@
The secondary constraints, when expressed through the metric components, are equivalent to the equations $G^{0\mu}=0$ since